12 moles each of Fe2O3 and H2 react according to the following reaction:
What is the maximum amount of iron produced from this reaction?
1
Expert's answer
2020-05-20T13:45:14-0400

Fe2O3 + 3H2=2Fe+3H2O

since there is less hydrogen than iron oxide, not all iron oxide will react, therefore we will assume that we are starting from the chemical amount of hydrogen.

3 — 12

2 — x

x=(12*2)/3=8

m(Fe)=8*56=448g